The mean price for used cars is $10,312. A manager of a Kansas City used car dealership reviewed a sample of 50 recent used car sales at the dealership in an attempt to determine whether the population mean price for used cars at this particular dealership differed from the national mean. The prices for the sample of 50 cars are contained in the Excel Online file below. Construct a spreadsheet to answer the following questions. Open spreadsheet a. Formulate the hypotheses that can be used to determine whether a difference exists in the mean price for used cars at the dealership. b. What are the test statistic and the p -value? Do not round intermediate calculations. t- (to 3 decimals) p-value (to 4 decimals) c. At a=0.01 what is your conclusion? the null hypothesis. We have evidence to conclude that the population mean price at this dealership differs from the national mean price of $10,312.

11611
7971
12838
11109
6474
6320
9523
6411
12805
9688
9478
10149
10176
11463
7084
8906
9508
9386
6506
8864
8447
7165
9591
7075
8431
6267
11309
8527
10488
9033
12180
6692
9622
7949
12730
9136
12588
12809
11024
11054
9743
10136
12629
8361
12188
11103
10578
7561
6623
8151
Answer #1

Given:

= 10.312, n = 50, = 9509.20, SD(S) = 2018.76, = 0.01

a)

Hypothesis:

Ho: = 10312

Ha: 10312

b)

Test statistic:

Degree of freedom = n-1 = 50-1 = 49

P-value = 0.0071        ....................Using t table

C)

Conclusion:

P-value < , i.e. 0.0071 < 0.05, That is Reject Ho at 1% level of signfiiacnce.

We have evidence to conclude that the population mean price at this dealership differs from the national mean price of $10,312.

We were unable to transcribe this image

We were unable to transcribe this image

We were unable to transcribe this image

We were unable to transcribe this image

We were unable to transcribe this image

We were unable to transcribe this image

X-4 t=slv 9509.20 - 10312 F= -2.812 2018.76/50

We were unable to transcribe this image